# CARZero: Cross-Attention Alignment for Radiology Zero-Shot Classification

### 27 Feb 2024 v1

@@ Abstract
The advancement of Zero-Shot Learning in the medical domain has been driven forward by using pre-trained models on large-scale image-text pairs, focusing on image-text alignment. However, existing methods primarily rely on cosine similarity for alignment, which may not fully capture the complex relationship between medical images and reports. To address this gap, we introduce a novel approach called Cross-Attention Alignment for Radiology Zero-Shot Classification (CARZero). Our approach innovatively leverages cross-attention mechanisms to process image and report features, creating a Similarity Representation that more accurately reflects the intricate relationships in medical semantics. This representation is then linearly projected to form an image-text similarity matrix for cross-modality alignment. Additionally, recognizing the pivotal role of prompt selection in zero-shot learning, CARZero incorporates a Large Language Model-based prompt alignment strategy. This strategy standardizes diverse diagnostic expressions into a unified format for both training and inference phases, overcoming the challenges of manual prompt design. Our approach is simple yet effective, demonstrating state-of-the-art performance in zero-shot classification on five official chest radiograph diagnostic test sets, including remarkable results on datasets with long-tail distributions of rare diseases. This achievement is attributed to our new image-text alignment strategy, which effectively addresses the complex relationship between medical images and reports. Code and models are available at TODO.
@@ 2.1 Zero-shot classification
For Vision Language Pretraining (VLP) tasks, previous works [li2019visualbert, li2021align] mainly use a fusion module to integrate image and text features, employing binary cross-entropy for classifying the combined features to determine if the image-text pair matches. Recently, CLIP [radford2021learning] introduced contrastive learning, which measures the cosine similarity between image and text features, aiming to maximize it between the matching image-text pairs and minimize the unpaired ones. This work significantly advances the development of VLP for ZSL in visual recognition tasks. Following CLIP, many studies [conde2021clip, ma2022ei, chen2023disco, dong2023maskclip, xie2023ra] have utilized contrastive learning for aligning image-text, demonstrating the substantial potential of contrastive learning for VLP.
@@ 3.1 Feature Extraction
Assume that the training dataset contains $N$ samples denoted as $D_{\text{train}}=\left\{\left(x_{1},y_{1}\right),\ldots,\left(x_{N},y_{N}\right)\right\}$, where $x_{i}\in\mathbb{R}^{H\times W\times 3}$ represents a CXR image and $y_{i}$ represents its corresponding medical report. $H$ and $W$ refer to height and width, respectively. As illustrated in Figure 2, we introduce individual components of our architecture for feature extraction of different modalities, including an image encoder $\Phi_{\text{image}}$ and a text encoder $\Phi_{\text{text}}$.
@@
Text Encoder As illustrated in Eq. 2, the text encoder is employed to extract text features from the given report. For each sample, a sentence is randomly selected from the report in each training iteration. In our experiments, to better adapt the text encoder for medical-related information extraction, we fine-tune PubMedBERT [gu2021domain] with the clinical reports and use it for text encoding.
@@ 3.2 Cross-Attention Alignment
Figure 3: Messages are used to prompt LLM for generating prompt templates from medical reports and diseases. Diseases are derived through keyword matching from reports related to 14 common diseases in the ChestXray14. Each prompt includes a manually curated one-shot example for enhanced specificity and relevance.
@@ 3.3 LLM-based Prompt Alignment
To align the prompts used during training and inference phases, we incorporate prompt templates into the training data using LLMs. As shown in Figure 3, prompting instruction is utilized to generate the prompt template within the training data. By leveraging the LLM’s exceptional capability for semantic understanding, fixed prompt templates are introduced into the training data. The templates generated by the LLM are then merged with the original reports to create an enhanced dataset for medical reporting. During the inference phase, the prompt template, “There is [disease].”, is employed for zero-shot classification. This strategy leverages the LLM’s advanced capabilities in semantic comprehension to ensure that the training data is enriched with consistent and relevant prompt templates, thereby facilitating a more effective and aligned application during both training and inference stages.
@@ 4.1 Dataset
MIMIC-CXR [johnson2019mimic] In our experiments, we conducted model pre-training using the MIMIC-CXR dataset, a publicly available collection of chest radiographs paired with radiology text reports. The MIMIC-CXR dataset comprises 377,110 images corresponding to 227,835 radiographic studies conducted on 65,379 patients. Each radiographic study is accompanied by a radiology report and the corresponding chest X-ray image, which may be in either frontal or lateral views. The radiology report serves as a comprehensive summary provided by radiologists, encompassing various sections such as examination, indication, impression, findings, technique, and comparison. In our methodology, we selectively retain only the findings and impressions sections from these reports. Moreover, only frontal views of CXRs are used for CARZero training.
@@
PadChest [bustos2020padchest] PadChest has 160,868 chest X-ray images labeled with 192 different diseases, which is a long-tailed distribution dataset. Only 27% of the labels, which equates to 39,053 examples, are attributed to board-certified radiologists. For evaluation purposes, we only test on samples annotated by board-certified radiologists. Additionally, we select categories with fewer than 10 samples, totaling 20 classes, designated as PadChest20, to evaluate the performance of CARZero on rare diseases.
@@ 4.3 Implementation Details
In our experiments, ViT-B/16 is used as the image encoder. For the text encoder, PubMedBERT is fine-tuned using texts from both MIMIC and PadChest datasets. Given that PadChest reports are in Spanish, they are translated into English. The LLM named Spark [^1] https://xinghuo.xfyun.cn/ is employed to intelligently insert prompt templates into the training dataset.

### 24 Mar 2024 v2

@@ Abstract
The advancement of Zero-Shot Learning in the medical domain has been driven forward by using pre-trained models on large-scale image-text pairs, focusing on image-text alignment. However, existing methods primarily rely on cosine similarity for alignment, which may not fully capture the complex relationship between medical images and reports. To address this gap, we introduce a novel approach called Cross-Attention Alignment for Radiology Zero-Shot Classification (CARZero). Our approach innovatively leverages cross-attention mechanisms to process image and report features, creating a Similarity Representation that more accurately reflects the intricate relationships in medical semantics. This representation is then linearly projected to form an image-text similarity matrix for cross-modality alignment. Additionally, recognizing the pivotal role of prompt selection in zero-shot learning, CARZero incorporates a Large Language Model-based prompt alignment strategy. This strategy standardizes diverse diagnostic expressions into a unified format for both training and inference phases, overcoming the challenges of manual prompt design. Our approach is simple yet effective, demonstrating state-of-the-art performance in zero-shot classification on five official chest radiograph diagnostic test sets, including remarkable results on datasets with long-tail distributions of rare diseases. This achievement is attributed to our new image-text alignment strategy, which effectively addresses the complex relationship between medical images and reports. Code and models are available at https://github.com/laihaoran/CARZero.
@@ 2.1 Zero-shot classification
For Vision Language Pretraining (VLP) tasks, previous works [li2019visualbert, li2021align] mainly use a fusion module to integrate image and text features, employing binary cross-entropy for classifying the combined features to determine if the image-text pair matches. Recently, CLIP [radford2021learning] introduced contrastive learning, which measures the cosine similarity between image and text features, aiming to maximize it between the matching image-text pairs and minimize the unpaired ones. This work significantly advances the development of VLP for ZSL in visual recognition tasks. Following CLIP, many studies [conde2021clip, ma2022ei, chen2023disco, dong2023maskclip, xie2023ra, varma2023villa, gandelsman2023interpreting, sun2023alpha] have utilized contrastive learning for aligning image-text, demonstrating the substantial potential of contrastive learning for VLP.
@@ 3.1 Feature Extraction
Assume that the training dataset contains $N$ samples denoted as $D_{\text{train}}=\left\{\left(x_{1},y_{1}\right),\ldots,\left(x_{N},y_{N}\right)\right\}$, where $x_{i}\in\mathbb{R}^{H\times W\times C}$ represents a CXR image and $y_{i}$ represents its corresponding medical report. $H$, $W$, and $C$ refer to height, width, and channels, respectively. As illustrated in Figure 2, we introduce individual components of our architecture for feature extraction of different modalities, including an image encoder $\Phi_{\text{image}}$ and a text encoder $\Phi_{\text{text}}$.
@@
Text Encoder As illustrated in Eq. 2, the text encoder is employed to extract text features from the given report. For each sample, a sentence is randomly selected from the report in each training iteration. In our experiments, to better adapt the text encoder for medical-related information extraction, we fine-tune BioBERT [lee2020biobert] with the clinical reports and use it for text encoding.
@@ 3.3 LLM-based Prompt Alignment
To align the prompts used during training and inference phases, we incorporate prompt templates into the training data using LLMs. Prompting instruction is utilized to generate the prompt template within the training data, with details shown in the supplementary. By leveraging the LLM’s exceptional capability for semantic understanding, fixed prompt templates are introduced into the training data. The templates generated by the LLM are then merged with the original reports to create enhanced reports for training. During the inference phase, the prompt template, “There is [disease].”, is employed for zero-shot classification. This strategy leverages the LLM’s advanced capabilities in semantic comprehension to ensure that the training data is enriched with consistent and relevant prompt templates, thereby facilitating a more effective and aligned application during both training and inference stages.
@@ 4.1 Dataset
MIMIC-CXR [johnson2019mimic] In our experiments, we conducted model pretraining using the MIMIC-CXR dataset, a publicly available collection of chest radiographs paired with radiology text reports. The MIMIC-CXR dataset comprises 377,110 images corresponding to 227,835 radiographic studies conducted on 65,379 patients. Each radiographic study is accompanied by a radiology report and the corresponding chest X-ray image, which may be in either frontal or lateral views. The radiology report serves as a comprehensive summary provided by radiologists, encompassing various sections such as examination, indication, impression, findings, technique, and comparison. In our methodology, we selectively retain only the findings and impressions sections from these reports. Moreover, only frontal views of CXRs are used for CARZero training.
@@
PadChest [bustos2020padchest] PadChest has 160,868 chest X-ray images labeled with 192 different diseases, which is a long-tailed distribution dataset. 39,053 (27%) samples are manually annotated by board-certified radiologists. For evaluation purposes, we only test on samples annotated by board-certified radiologists. Additionally, we select categories with fewer than 10 samples, totaling 20 classes, designated as PadChest20, to evaluate the performance of CARZero on rare diseases.
@@ 4.2 Evaluation Metric
MedCLIP [wang2022medclip] 0.551 0.508 0.501 0.564 0.744 0.571
BiomedCLIP [zhang2023large] 0.577 0.513 0.510 0.639 0.677 0.630
@@ 4.3 Implementation Details
In our experiments, ViT-B/16 is used as the image encoder, which utilizes M3AE [chen2022multi] for pretraining on the MIMIC dataset. For the text encoder, BioBERT is fine-tuned using texts from both MIMIC and PadChest datasets. Given that PadChest reports are in Spanish, they are translated into English. The LLM named Spark [^1] https://xinghuo.xfyun.cn/ is employed to intelligently insert prompt templates into the training dataset.
@@ 5 Conclusion, Limitation and Impact
6 Acknowledgment
This work was supported by Natural Science Foundation of China under Grant 62271465 and Open Fund Project of Guangdong Academy of Medical Sciences, China (No. YKY-KF202206).